35. Find the sum of this series: –1458 + 486 – 162 + 54 – … ; 7th term.
Answer by htmentor(1343)

The sum of n terms of a geometric series is S_n = a*(1-r^n)/(1-r) where a is the 1st term and r is the common ratio.
This is a geometric series with r = -1/3
Thus S_7 = -1458*(1-(-1/3)^7)/(1-(-1/3)) = -1094